OpenStudy (anonymous):

How do I simplify: (x+1/2) x (x-3/2)?

OpenStudy (saifoo.khan):

we will use FOIL.

OpenStudy (saifoo.khan):

x^2 - 3x/2 + 1x/2 - 3/4

OpenStudy (saifoo.khan):

Add like terms, x^2 -x -3/4

OpenStudy (anonymous):

How do I simplify an expression that represents the fraction of the area of the large rectangle, which is what I placed above, covered by the shaded rectangle: l=(x+1/4), w= x-2

OpenStudy (saifoo.khan):

same way.

OpenStudy (saifoo.khan):

we will multiply both parenthesis.